The plane x + 2y – z = 4 cuts the sphere x2 + y2 + z2 – x + z – 2 = 0 in a circle of radius ?

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Perpendicular distance of centre(1/2,0,-1/2)from x + 2y – 2 = 4
=(|1/2+1/2-4|)/\sqrt{6}=\sqrt{3/2}
So radius=\sqrt{5/2-3/2}=1
Thanks and Regards
Shaik Aasif
askIITians faculty